OpenStudy (anonymous):

HELP PLEASE! simplify the radical expressions. state any excluded values. show your work. 1. 3x-6/x-2 2. x-2/x^2+3x-10

OpenStudy (anonymous):

combined like turms and bring all the x's together then combined the regular numbers

OpenStudy (anonymous):

i've tried. i cant understand how to do it.. i suck at math..

OpenStudy (akashdeepdeb):

\[1.~~\frac{3x-6}{x-2} = \frac{3(x-2)}{(x-2)} = \frac{3 \cancel{(x-2)}}{\cancel{(x-2)}} = 3\] \[2. ~~\frac{x-2}{x^2 + 3x - 10} = \frac{(x-2)}{(x+5)(x-2)} =\frac{\cancel{(x-2)}}{(x+5)\cancel{(x-2)}} = \frac{1}{(x+5)}\] Getting this? :)

OpenStudy (anonymous):

getting it a little better. i suck at math bad..

OpenStudy (anonymous):

3x goes with x so 4 x and 2 negitives equals a positive so -6 * -2 = 12 12/4x equals x =3

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

yeah 3 is the answer

OpenStudy (anonymous):

so would x-2 be the excluded values? or would it be 3?